OpenStudy (anonymous):

Find the limit of the sequence an= ln(9n+7) - ln(n)

OpenStudy (anonymous):

I said it would Diverge, but it's incorrect hahaha

OpenStudy (anonymous):

does it help to know that \[\ln(9n+7)+\ln(n)=\ln\left(\frac{9n+7}{n}\right)\]?

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

negative that's okay

OpenStudy (anonymous):

\[\ln(9n+7)-\ln(n)=\ln\left(\frac{9n+7}{n}\right)\]

OpenStudy (anonymous):

converge to ln (9)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you very much :)

OpenStudy (freckles):

I would say 9n+7 is approximately 9n for large n ln(9n/n)=ln(9) :)

OpenStudy (anonymous):

Good point there